The product of two consecutive integers is 8 less than 8 times their sum. Find the integers
Answer by JimboP1977(311)

x*(x+1)-8 = 8(x+x+1)
x^2+x-8= 16x+8
x^2-15x=0
(x-15)(x+0)=0
So x = 15
So two numbers are 15 and 16.